# Gravity-Informed Deep Learning Framework for Predicting Ship Traffic Flow and Invasion Risk of Non-Indigenous Species via Ballast Water Discharge

### v1

@@ Acknowledgments
This research was partially funded by the Natural Sciences and Engineering Research Council of Canada (NSERC), Grant number RGPIN-2022-03909; Sao Paulo State Research Foundation (FAPESP), Grant number 2016/17078-0 and 2022/12374-1; the Institute for Big Data Analytics (IBDA) and the Ocean Frontier Institute (OFI) at Dalhousie University, Halifax - NS, Canada; and further funded by the Canada First Research Excellence Fund (CFREF), the Canadian Foundation for Innovation MERIDIAN cyberinfrastructure. We thank all funding sources and Spire for providing the vessel trajectory dataset.
@@
Conceptualization, R.S. and G.S.; methodology, R.S., G.S. and A.S.; validation, R.S., G.S., and A.S.; formal analysis, R.S., G.S., and S.B.; data modeling, R.S. and G.S.; data curation, R.S., G.S., and A.S.; writing — original draft preparation, R.S. and G.S.; writing — review and editing, all authors; visualization, R.S.; supervision, R.P., S.M., and A.S.; project administration, G.S. and A.S.; funding acquisition, R.P., S.M., A.S.; All authors have read and agreed to the published version of the manuscript.

### v3

@@ Acknowledgments
This research was partially funded by the Natural Sciences and Engineering Research Council of Canada (NSERC), Grant number RGPIN-2022-03909; Sao Paulo State Research Foundation (FAPESP), Grant number 2016/17078-0 and 2022/12374-1; the Institute for Big Data Analytics (IBDA) and the Ocean Frontier Institute (OFI) at Dalhousie University, Halifax - NS, Canada; and further funded by the Canada First Research Excellence Fund (CFREF), the Canadian Foundation for Innovation MERIDIAN cyberinfrastructure. We thank all funding sources and Spire for providing the vessel trajectory dataset. We also thank Sarah Bailey, from the Great Lakes Laboratory for Fisheries and Aquatic Sciences, Fisheries and Oceans, Dartmouth – NS, Canada, for reviewing and sharing her thoughts about this paper.
@@
Conceptualization, R.S. and G.S.; methodology, R.S., G.S. and A.S.; validation, R.S., G.S., and A.S.; formal analysis, R.S., and G.S.; data modeling, R.S. and G.S.; data curation, R.S., G.S., and A.S.; writing — original draft preparation, R.S. and G.S.; writing — review and editing, all authors; visualization, R.S.; supervision, G.S., A.S, R.P., and S.M.; project administration, G.S. and A.S.; funding acquisition, A.S., S.M., R.P., and G.S.; All authors have read and agreed to the published version of the manuscript.
Data Availability
The research we describe in this paper has used data acquired from Spire under MERIDIAN’s fair-use and non-disclose data license, which is research-intended high-resolution data that is hosted at the Institute for Big Data Analytics (IBDA) from Dalhousie University. Due to licensing agreements, we cannot share the raw data, but we can provide a trained model instead; contact the corresponding author for assistance.